### v1

# Bounds for Lyapunov exponent of circular light orbits in black holes

Abstract
Chaotic systems near black holes satisfy a universal bound, $\lambda\leq\kappa_{H}$ linking the Lyapunov coefficient $\lambda$ associated with unstable orbits to surface gravity $\kappa_{H}$ of the event horizon. A natural question is whether this bound is satisfied by unstable circular null geodesics in the vicinity of black holes. However, there are known cases where this bound is violated. It is intriguing to ask whether there exists an alternative universal bound that is valid in such situations. We show that for any spherically symmetric, static black hole that satisfies Einstein’s equations and the dominant energy condition, there exist other universal bounds relating the Lyapunov coefficient to a generalized notion of surface gravity at the photon sphere, as well as to the Unruh temperature locally measured by static observers. As applications, we show how these bounds also constrain the imaginary part of quasinormal modes in the eikonal regime and how the Lyapunov coefficient relates to the shadow size and the entropy of the horizon.
I Introduction
Lyapunov exponents, $\lambda$, quantify how rapidly initially close trajectories in phase space either converge or diverge from one another. A positive Lyapunov exponent indicates that nearby trajectories will diverge quickly, reflecting a high sensitivity to initial conditions. In their seminal paper Maldacena et al. (2016), Maldacena and collaborators identified a universal upper bound on chaos in a thermal quantum field theory at temperature $\tilde{T}$. This bound is characterized by the Lyapunov exponent of out-of-time-ordered correlators and has in the natural units the simple expression $\lambda\leq 2\pi\tilde{T}$. Further investigations uncovered the validity of this universal bound in other systems, like for massive particle motions near black hole horizons Hashimoto and Tanahashi (2017), a connection to the presence of a horizon Dalui et al. (2019), and how potential quantum corrections might enhance chaotic behavior of massless particles near the horizon Bera et al. (2022). Indeed, many of the investigated systems satisfy the inequality
$\lambda\leq\kappa_{H}={2\pi T_{H}},$ (1)
where $\kappa_{H}$ is the surface gravity of the horizon of the black hole and $T_{H}$ its associated temperature.
Circular orbits of null geodesics play a role in upcoming space-based very long baseline interferometry missions. The flux ratio between consecutive light rings can be related to the Lyapunov exponents of corresponding adjacent nearly bound null geodesics Johnson et al. (2020). Thus an insight into the Lyapunov exponents of these orbits offers a better understanding of black hole observations and thereby provides means to test modified theories of gravity. In fact, several authors analyzed such dependence for various spacetime geometries with the aim of placing constraints on alternative theories through long-baseline observations Staelens et al. (2023); Deich et al. (2024); Salehi et al. (2024); Gralla and Lupsasca (2020); Kocherlakota et al. (2024); Wong et al. (2024).
The Lyapunov exponent associated with unstable circular-photon null geodesics is also related to lensing in the strong deflection limit of black holes Stefanov et al. (2010), to the imaginary part of quasinormal modes in the eikonal limit Cardoso et al. (2009), and to the phase-transition properties of black holes which multiple branches correspond to the distinct phases of the black hole Guo et al. (2022); Kumara et al. (2024); Shukla et al. (2024).
From the observational importance of the Lyapunov exponent arises the natural question: Does relation (1) hold for unstable null circular geodesics? As a matter of fact for many spacetimes the answer is positive, but a corresponding relation for near-extremal black holes does not hold Yu et al. (2022); Bianchi et al. (2020). The reason for the latter is that in the limit towards extremality the surface gravity $\kappa_{H}$ approaches zero, but the Lyapunov exponent does not. Consequently, there is a region in the parameter space of nearly extremal black holes where the bound is violated.
Nevertheless, Abreu and Visser (2010) introduced a concept of a generalized surface gravity, $\kappa(r)$, for static, spherically symmetric spacetimes, which is applicable at arbitrary radial positions, $r$. In particular, evaluation of $\kappa(r)$ at the position $r_{H}$ of a black hole horizon yields its surface gravity $\kappa(r_{H})=\kappa_{H}$. Moreover, at the radial position $r_{\gamma}$ of unstable circular null geodesics, it is non-vanishing even for nearly extremal black holes. This provokes the question if a similar relation to eq. (1) holds when replacing $\kappa_{H}$ by the generalized surface gravity $\kappa_{\gamma}\equiv\kappa(r_{\gamma})$. That is, setting a bound to the Lyapunov exponent of circular null geodesics with the value of the corresponding generalized surface gravity at $r_{\gamma}$.
Our answer to this question is affirmative. For any spherically symmetric black hole satisfying the Einstein equations and the dominant energy condition, there exists a universal upper bound relating $\lambda$ and $\kappa_{\gamma}$. This upper bound also explicitly depends on the radial position of the photon sphere $r_{\gamma}$. Furthermore, we show the existence of a bound between the Lyapunov exponent and the Unruh temperature $T_{U}$ measured by static observers located at $r_{\gamma}$. This bounding relation, $\lambda\leq 2\pi T_{U}$, is saturated for a Schwarzschild black hole. In fact, it was recently derived in Giataganas et al. (2024) for the two particular cases of Schwarzschild and Reissner-Nordström black holes. Our proof generalizes their results to a broader class of spherically symmetric black holes and is sufficiently general to not require an explicit form of the metric. Moreover, our new bound is directly related to recent bounds associated with the shadow size and acceleration bounds for radial linear uniformly accelerated trajectories in static spherically symmetric black hole spacetimes of the Schwarzschild type Paithankar and Kolekar (2023).
Our major results culminate in the expressions for the generic bounds between the Lyapunov exponent and (i) surface gravity at the photons sphere (eq. (38), (39) and (40)), (ii) the Unruh temperature of static observers at the photon sphere (eq. (42)), and (iii) the shadow of the black hole (eq. (44)).
As an application of these new bounds, we derive a constraint for the quasinormal modes. Following the approach in Giataganas et al. (2024), we also present simple bounds for the Lyapunov exponent using the inverse of the entropy of the black hole horizon.
The note is organized as follows: in Sec. II we outline the general geometric setup; in Sec. III, we derived the different expression for the bounds of the Lyapunov exponents; and in Sec. IV, the major results as summarized and possible future work is indicated. Natural units $G=c=h=1$ are used throughout and the metric signature is $+2$.
II Background
Consider a 4-dimensional, spherically symmetric and static spacetime with a metric described by
$\begin{split}ds^{2}=&-\frac{\mu(r)}{e^{2\delta(r)}}dt^{2}+\frac{dr^{2}}{\mu(r)}+r^{2}\left(d\theta^{2}+\sin^{2}\theta\,d\phi^{2}\right),\end{split}$ (2)
We assume that the spacetime contains an event horizon of a black hole at $r=r_{H}>0$. We have $\mu(r_{H})=0$ at the event horizon and require the additional conditions Nunez et al. (1996)
$\mu^{\prime}(r_{H})\geq 0\;\;,\delta(r_{H})\leq\infty\;\;,\;\;\delta^{\prime}(r_{H})\leq\infty\;\;.\;\;$ (3)
In the exterior $r>r_{H}$, we demand the asymptotical flatness conditions
$\lim_{r\rightarrow\infty}\mu=1\;\;,\;\;\lim_{r\rightarrow\infty}\delta=0\;\;.\;\;$ (4)
Einstein equations are $G_{ab}=8\pi T_{ab}$ with the Einstein tensor $G_{ab}$ and energy-momentum tensor $T_{ab}$ with corresponding trace $T:=g^{ab}T_{ab}$. The energy density as measured by static observers is given by $\rho=-T_{ab}t^{a}t^{b}$ with the timelike unit vector $t^{a}$ which is proportional to the timelike Killing vector field $\partial_{t}$. Based on $t^{a}$ we form the orthonormal tetrad ${E_{\alpha}=(t^{a},e^{a}_{(r)},e^{a}_{(\theta)},e^{a}_{(\phi)})}$. This tetrad defines the radial pressure $p_{r}=T_{ab}e^{a}_{(r)}e^{b}_{(r)}$, and the pressure $p_{\bot}$ tangent to the orbits of spherical symmetry, i.e. $p_{\bot}=T_{ab}e^{a}_{(\theta)}e^{b}_{(\theta)}=T_{ab}e^{a}_{(\phi)}e^{b}_{(\phi)}$. Consequently, the trace of the energy momentum tenor is ${T=-\rho+p_{r}+2p_{\bot}}$. We require the dominant energy condition Hawking and Ellis (2023); Wald (1984), $\rho\geq 0$, $\rho\geq|p_{r}|$ and $\rho\geq|p_{\bot}|$.
The function $\mu$ relates to the Misner-Sharp mass $m(r)$ Misner and Sharp (1964)
[…]
where we defined the horizon mass $m_{H}$. The ADM mass $M$ is the obtained from the limit of $m$ as $r\to\infty$. Finiteness of the limit of $m$ implies by (5) that
$\lim_{r\rightarrow\infty}r^{3}\rho=0\;\;.$ (6)
From the $(tt)$ and $(rr)$ components of field equations and from $\nabla_{a}T^{a}_{r}=0$ we get (Gallo and Villanueva, 2015, for $\hat{\alpha}=0$ and $n=4$)
$8\pi\rho=$ $=\frac{1}{r^{2}}-\frac{(r\mu)^{\prime}}{r^{2}}\;,$ (7)
$8\pi p_{r}=$ $\frac{e^{2\delta}}{r^{2}}\left(\frac{r\mu}{e^{2\delta}}\right)^{\prime}-\frac{1}{r^{2}}\;,$ (8)
$p^{\prime}_{r}=$ $-\frac{(e^{-2\delta}\mu)^{\prime}}{2e^{-2\delta}\mu}(\rho+p_{r})+\frac{2}{r}(p_{\bot}-p_{r})\;.w$ (9)
where denotes the derivative with respect to $r$. The sum of (7) and (8) and evaluation at the horizon gives with $\mu(r_{H})=0$ that $\rho(r_{H})+p_{r}(r_{H})=0$. Based on eqs.(7) and (8), we have
$\mu\delta^{\prime}=-4\pi r(\rho+p_{r}).$ (10)
For any metric of the form (2), the radial coordinate of null geodesics satisfies an equation of the form $\dot{r}^{2}=V(r,E,L),$ where a dot denotes a derivative with respect to an affine parameter, and $V$ is an effective potential given by Cardoso et al. (2009)
$V=e^{2\delta(r)}E^{2}-\frac{L^{2}}{r^{2}}\mu(r),$ (11)
with $E$ the energy and $L$ the orbital angular momentum of the massless particle. Circular null geodesics at ${r=r_{\gamma}}$ satisfy $V=V^{\prime}=0$. For the metric (2), it is equivalently to require $\tilde{N}(r_{\gamma})=0$ (see for example Cardoso et al. (2009); Hod (2013)) with
$\tilde{N}(r)=2e^{-2\delta(r)}\mu(r)-r[e^{-2\delta(r)}\mu(r)]^{\prime}.$ (12)
Alternatively, circular null geodesics can be obtained from the vanishing of the function $N(r)=e^{2\delta(r)}\tilde{N}(r)$, which by using (8) can be expressed as
$N(r)=3\mu(r)-1-8\pi r^{2}p_{r}.$ (13)
As first shown by Hod Hod (2013), eq. (13) admits at least one solution, which follows from the following facts: First, the conditions $\mu(r_{H})=0$ and (3) imply that $N(r)$ satisfies $N(r_{H})\leq 0$. Second, taking into account that $\lim\limits_{r\rightarrow\infty}r^{2}p_{r}=0$, which follows from the dominant energy condition and (6), we see that $N(r\to\infty)\to 2$. Therefore, there exists a value $r_{\gamma}$ such that $N(r_{\gamma})=0$. Since we are interested in the innermost null circular orbit defined by $r=r_{\gamma}$, $N(r)$ satisfies
$N(r_{H}\leq r<r_{\gamma})<0,$ (14)
and
$N^{\prime}(r_{\gamma})\geq 0.$ (15)
The stability of the orbits follows from considering the second derivative of the effective potential $V$ given by (11), whose evaluation at $r_{\gamma}$ gives
$V^{\prime\prime}({r_{\gamma}})=\frac{L^{2}}{r^{4}e^{-2\delta{(r)}}}\left[2e^{-2\delta{(r)}}\mu(r)-r^{2}_{\gamma}[e^{-2\delta(r)}\mu(r)]^{\prime\prime}\right]\bigg{|}_{r_{\gamma}},$ (16)
which, using the Einstein equations (7)-(9), can be rewritten as Gallo and Villanueva (2015),
$V^{\prime\prime}(r_{\gamma})=\frac{L^{2}N^{\prime}(r_{\gamma})}{r_{\gamma}^{3}}.$ (17)
Therefore, taking into account (15), we obtain $V^{\prime\prime}(r_{\gamma})\geq 0$, implying that in general $r_{\gamma}$ corresponds to an unstable circular orbit [^1] Note that the equation $N(r)=0$ could have more than one solution. From its behavior at $r_{H}$ and at the asymptotic region the outer light ring must be unstable. However, here we are interested in the inner circular orbit where $N^{\prime}(r_{\gamma})\geq 0$. If $N^{\prime}(r_{\gamma})>0$, the instability follows. But for the special case of $N^{\prime}(r_{\gamma})=0$, the stability behavior would depend on the behavior of higher derivatives of $N(r)$. However, the case $N^{\prime}(r_{\gamma})=0$ is only possible for a very special case where the matter content is such that $(\rho(r_{\gamma})+p_{\bot}(r_{\gamma}))=(8\pi r^{2}_{\gamma})^{-1}$ (see Eq.(37)). .
Since most matter fields obey $T\leq 0$, Hod Hod (2013) used this condition to prove that $\mu(r_{\gamma})$ obeys
$\mu(r_{\gamma})\leq\frac{1}{3},$ (18)
which by using (5) gives
$r_{\gamma}\leq 3{m(r_{\gamma})}\leq 3{M}.$ (19)
III Bounds on the Lyapunov exponent of circular null geodesics
For massive particle motion near the horizon of a black hole, there exists a relation between the Lyapunov exponent $\lambda$ of their phase space trajectories and the surface gravity $\kappa_{H}$ of the event horizon, specifically, $\lambda\leq\kappa_{H}.$
Hereafter, we briefly summarize the computation of the Lyapunov exponents for the motion of test particles, following Cornish and Levin (2003); Cardoso et al. (2009). The equations of motion of $N$ particles can be schematically be written as
$\frac{dX_{i}}{dt}=H_{i}(X_{j},g_{ab}),$ (20)
where $X_{i}$ is a 2N dimensional vector consisting of the position and velocities (or momenta) of the particles and $H_{i}(X_{j})$ is a $2N$ dimensional vector depending on $X_{i}$, the metric and the metric’s first derivatives. Linearisation around an orbit $X_{i}$ yields
$\frac{d\delta X_{i}(t)}{dt}=K_{ij}(t)\delta X_{j}(t),$ (21)
where $K_{ij}(t)=\frac{\partial H_{i}}{\partial X_{j}}\bigg{|}_{X_{i}(t)}$ is the stability matrix. A solution to (21) is
$\delta X_{i}(t)=L_{ij}(t)\delta X_{j}(0),$ (22)
with $\dot{L}_{ij}(t)=K_{im}L_{mj}(t)$ and $L_{ij}(0)=\delta_{ij}$.
The Lyapunov exponent $\lambda$ is determined by
$\lambda=\lim_{t\to\infty}\frac{1}{t}\log\left[\frac{L_{jj}(t)}{L_{jj}(0)}\right].$ (23)
For one particle with a two-dimensional phase space $X_{i}(t)=(p_{r},r)$, such as circular orbits in spherically symmetric spacetimes, the equations linearize to
$K_{ij}=\begin{pmatrix}0&K_{1}\\ K_{2}&0\end{pmatrix}.$ (24)
with stability matrix components given by
$K_{1}$ $=$ $\frac{d}{dr}\left(\dot{t}^{-1}\frac{\partial\mathcal{L}}{\partial r}\right),$ (25)
$K_{2}$ $=$ $-\dot{t}g_{rr}^{-1},$ (26)
where $\mathcal{L}=\frac{1}{2}g_{\alpha\beta}\dot{x}^{\alpha}\dot{x}^{\beta}$ the Lagrangian for geodesic motion. The principal Lyapunov exponents for circular orbits can then be expressed as
$\lambda^{2}=K_{1}K_{2}.$ (27)
In particular, the Lyapunov exponent for circular null geodesics of metrics of the form (2) is given by Cardoso et al. (2009)
$\lambda=\sqrt{\frac{r_{\gamma}^{2}[e^{-2\delta(r_{\gamma})}\mu(r_{\gamma})]}{2L^{2}}V^{\prime\prime}(r_{\gamma})},$ (28)
which while using (17) yields Gallo and Villanueva (2015)
$\lambda=\sqrt{\frac{e^{-2\delta}\mu}{2r}N^{\prime}}\bigg{|}_{r_{\gamma}},$ (29)
which is a real quantity due to eq. (15).
For null geodesics at the photon sphere, the bound (1) can be violated where $\lambda$ is computed via (29). Examples are near-extremal Reissner-Nordström black holes with a charge $Q\in[0.99M,M]$ as shown in Bianchi et al. (2020).
To establish a bound relating the Lyapunov exponent with a surface gravity, which does not exclude these cases, we employ the notion of the generalized surface gravity $\kappa(r)$ of Abreu and Visser (2010) which is valid for every sphere of radius $r$. The usage of $\kappa(r)$ provides simple bounds on quasi-local entropies similar to the holographic bounds relating entropy with area of the enclosed region. This generalized surface gravity $\kappa(r)$ is determined by a redshift factor of the norm $a=(g_{bc}a^{b}a^{c})^{1/2}$ of the acceleration $a^{b}=t^{c}\nabla_{c}t^{b}$ of observers following the integral curves along the unit vector field $t^{c}$. For (2), it results in
$\kappa(r)=\sqrt{e^{-2\delta}\mu}a(r),$ (30)
with
$a(r)=\frac{m(r)-rm^{\prime}(r)}{r^{2}\sqrt{\mu}}-{\delta^{\prime}(r)}{\sqrt{\mu}}.$ (31)
Using the Einstein equations in (31), we obtain
$a(r)=\frac{1}{\sqrt{\mu}}\left(\frac{m(r)}{r^{2}}+4\pi rp_{r}\right),$ (32)
which allows us to rewrite $\kappa(r)$ as
$\kappa(r)=\frac{e^{-\delta}}{r^{2}}\left[m(r)+4\pi r^{3}p_{r}\right]=\frac{e^{-\delta}}{r}\left(\mu-\frac{N}{2}\right),$ (33)
where we have expressed $m(r)$ in terms of $\mu$, and $p_{r}$ in terms of $N$ using Eq.(13). In most physical situations, one would expect $\kappa(r)<\kappa_{H}$ in the exterior region of a black hole horizon Abreu and Visser (2010). However, for near-extremal black holes, where $\kappa_{H}\to 0$, the generalized surface gravity does remains positive for $r\rightarrow r_{\gamma}$, meaning $\kappa_{\gamma}=\kappa(r_{\gamma})$ does not vanish. Therefore, it is natural to seek an upper bound for the Lyapunov exponent in terms of $\kappa_{\gamma}$ instead of $\kappa_{H}$.
Evaluation of (33) at the photon radius $r_{\gamma}$ and using $N(r_{\gamma})=0$, we find
$\kappa_{\gamma}=\frac{e^{-\delta}\mu}{r}\bigg{|}_{r_{\gamma}}=\sqrt{e^{-2\delta(r_{\gamma})}{\mu(r_{\gamma})}}a_{U},$ (34)
where $a_{U}=a(r_{\gamma})$. This norm of the acceleration at $r_{\gamma}$ relates to the Unruh temperature $T_{U}(r_{\gamma})$, which is minimum local temperature of vacuum measured by static observers at $r_{\gamma}$
$a_{U}=\frac{\sqrt{\mu}}{r}\bigg{|}_{r_{\gamma}}=2\pi T_{U}(r_{\gamma}).$ (35)
To relate (29) with (33), we need to calculate the derivative of $N$ which will contain the derivative of $p_{r}$. For the latter, it follows from the vanishing of Eq.(12) and Eq.(9) that
$p^{\prime}_{r}(r_{\gamma})=\frac{1}{r}(2p_{\bot}-3\rho+p_{r})\bigg{|}_{r_{\gamma}}.$ (36)
Thus, by differentiating Eq.(13) with respect to $r$ and taking into account Eq.(7) and (36) together with the dominant energy condition, i.e. $\rho+p_{\bot}\geq 0$, we find at $r_{\gamma}$
$N^{\prime}(r_{\gamma})=\frac{2}{r}\left[1-8\pi r^{2}(\rho+p_{\bot})\right]\bigg{|}_{r_{\gamma}}\leq\frac{2}{r_{\gamma}}.$ (37)
Therefore, we are in condition to prove the following inequality:
$\begin{split}\lambda=&\sqrt{\frac{e^{-2\delta}\mu}{2r}N^{\prime}}\bigg{|}_{r_{\gamma}}\\ =&\frac{\kappa_{\gamma}}{\sqrt{\mu}_{\gamma}}\left({1-8\pi r^{2}(\rho_{\gamma}+p_{\bot\gamma})}\right)^{1/2}\leq\frac{\kappa_{\gamma}}{\sqrt{\mu_{\gamma}}},\end{split}$ (38)
relating $\lambda$ to $\kappa_{\gamma}$, but also depending on $r_{\gamma}$ through $\mu_{\gamma}$. Since in general at $r_{\gamma}$ we only know that $\mu_{\gamma}\leq\frac{1}{3}$ (assuming a no positive trace of $T^{ab}$), we see that the inequality $\lambda\leq\kappa_{\gamma}$ cannot be guaranteed by eq.(38). In fact, for an extremal Reissner-Nordström metric, we have $\lambda=\frac{\sqrt{2}}{8}$, while $\kappa_{\gamma}=\frac{1}{8}$, and therefore, $\lambda>\kappa_{\gamma}$. Nevertheless, in consistence with our proof, it can be checked that the bound in eq. (38) is satisfied for any Reissner-Nordström black hole, and in particular for an extremal Reissner-Nordström black hole where at $r_{\gamma}$, $\mu_{\gamma}=\frac{1}{4}$ [^2] For a RN black hole with ADM mass $M$ and charge $Q$, $m(r)=M-\frac{Q^{2}}{2r}$ and the outer photon sphere is placed at $r_{\gamma}=\frac{3}{2}M+\frac{1}{2}\sqrt{9M^{2}-8Q^{2}}$. In particular for an extremal Reissner-Nordström black hole, $r_{\gamma}=2M$, and $m(r_{\gamma})=3M/4$..
Alternatively, we can express (38) explicitly in terms of $\kappa_{\gamma}$ and $r_{\gamma}$. This follows from (4) and (10) and the dominant energy condition, which imply $\delta^{\prime}\leq 0$ and therfore $e^{-\delta(r_{\gamma})}\leq 1$. Under these circumstances, we get from (29) and (33)
$\begin{split}\lambda&=\sqrt{\frac{e^{-2\delta}\mu}{2r}N^{\prime}}\bigg{|}_{r_{\gamma}}=e^{-\delta}\sqrt{\frac{e^{-\delta}\mu}{r}\frac{{1-8\pi r^{2}(\rho_{\gamma}+p_{\bot\gamma})}}{r}}\bigg{|}_{r_{\gamma}}\\ &\leq e^{-\delta}\sqrt{\frac{\kappa}{r}}\bigg{|}_{r_{\gamma}}\leq\sqrt{\frac{\kappa_{\gamma}}{r_{\gamma}}}.\end{split}$ (39)
The bound (39) also implies $\lambda^{2}r_{\gamma}\leq{\kappa_{\gamma}}$. Since $r_{\gamma}>r_{H}$, we obtain the simple inequality
$\lambda^{2}r_{H}\leq\kappa_{\gamma}.$ (40)
It results that $\lambda$ is bounded by $\kappa_{\gamma}$ and for the size of the event horizon. In fact, this inequality can be improved for black holes with a traceless energy-momentum tensor. In such cases, Hod recently showed in Hod (2023) that $r_{\gamma}\geq\frac{6}{5}r_{H}$, which implies that $\lambda^{2}r_{H}\leq\frac{5}{6}\kappa_{\gamma}$.
It turns out that the relations (38), (39) and (40) are not the only inequalities which can be written at the photon sphere of spherically symmetric black hole under the given physical assumption. In fact, we can re-express the inequality (38) in a very natural way in terms of the Unruh temperature as measured by static observers placed at $r_{\gamma}$. To do so, note that from eq. (34) and the dominant energy condition, it follows that
$\frac{\kappa}{\sqrt{\mu}}\bigg{|}_{r_{\gamma}}=e^{-\delta(r_{\gamma})}a_{{U}}\leq a_{{U}},$ (41)
which by using (35) and (38) implies the universal upper bound for the Lyapunov exponent:
$\lambda\leq 2\pi T_{U}.$ (42)
It is worth noting that, recently, in reference Giataganas et al. (2024), the validity of this inequality was observed for the special cases of Schwarzschild and Reissner-Nordström black holes through direct calculation. What we have shown in Eq.(42) is that it remains valid within Einstein’s theory of general relativity, under very natural physical assumptions. In Giataganas et al. (2024), the authors also conjecture new inequalities that have been tested to hold for rotating spacetimes in the Kerr-Newman family (more on this will be discussed below.)
In turn, Paithankar and Kolekar (2023) showed that metrics like (2) but with $\delta(r)=0$, there exists an equality between the size of the black hole shadow and the acceleration $a_{b}$ associated with uniformly accelerated radial motion, which passes through $r_{\gamma}$. Specifically, they found that the radius of the shadow $R_{sh}$ is given by $R_{sh}=a_{b}^{-1}$. Moreover, it follows that $a_{b}=a_{U}$ so that $R_{sh}=a_{U}^{-1}$. As the most general radius $R_{sh}$ of the shadow is given by
$R_{sh}=\frac{r_{\gamma}}{\sqrt{e^{-2\delta(r_{\gamma})}\mu(r_{\gamma})}}=\frac{e^{\delta(r_{\gamma})}}{a_{U}}.$ (43)
We can see that we have a relation between the Unruh acceleration and the shadow radius, namely $a_{U}=e^{\delta(r_{\gamma})}R_{sh}^{-1}$, which provides a link between the Lyapunov exponent and the shadow size of the black hole using (38), (41) and (43),
$\lambda\leq\frac{1}{R_{sh}}.$ (44)
This inequality has already appeared in Bianchi et al. (2020) for the special family of spherically symmetric and charged Anti-de-Sitter metrics, where it was verified to be valid by explicit computation. In that reference, the inequality is expressed in terms of the impact parameter $b$ (which for an asymptotic observer coincides with $R_{sh}$). Once again, our results generalize this inequality for a broader family of black holes.
As final remarks, the bounds (38), (39), and (40) together with the assumption of a non-positive trace of the energy momentum tensor imply new (weaker) bounds relating $\lambda$ directly with $r_{\gamma}$ or $r_{H}$. Indeed, (18), (34), the dominant energy condition and the Einstein equations imply that ${\kappa_{\gamma}r_{\gamma}\leq\frac{1}{3}}$. Hence, as $r_{\gamma}>r_{H}$ we have from (40)
$\lambda\leq\frac{1}{\sqrt{3}r_{H}}.$ (45)
This inequality allows to write a universal bound for the imaginary part of the quasinormal modes in the eikonal limit. In that limit, defining $\omega=\omega_{R}-i\omega_{I}$ as the frequency of the quasinormal modes, we have $\omega_{I}=(n+\frac{1}{2})\lambda$ with $n$ the overtone numberCardoso et al. (2009). Therefore, we obtain the universal upper bound
$\omega_{I}r_{H}\leq\frac{1}{3\sqrt{3}}\left(n+\frac{1}{2}\right).$ (46)
An associated bound for the real $\omega_{R}$ was derived in Hod (2013).
The bound (45) can be related to a conjecture of Giataganas et al. (2024) relating the Lyapunov exponent to the black hole entropy $S$, which reads Giataganas et al. (2024)
$\lambda^{2}\leq\frac{\pi}{4S}.$ (47)
where $S$ is obtained from the area $A$ of the event horizon via $S=\frac{A}{4}=\pi r_{H}^{2}$.
In contrast, from Eq.(45), we obtain
$\lambda^{2}\leq\frac{1}{3r^{2}_{H}}=\frac{\pi}{3S},$ (48)
which is a weaker bound than (47). The inequality (48) holds for a broader family of spherically symmetric black holes than (47) of Giataganas et al. (2024). Their inequality was shown to remain valid for the Kerr-Newman family of black holes. Since (48) is a weaker bound, it is clear that it remains valid for the spacetimes studied in Giataganas et al. (2024). Naturally, for spherically symmetric black holes, writing the upper bound for $\lambda$ in terms of $r_{H}$ or $S$ is entirely equivalent. However, this equivalence does not hold for spinning black holes. It would be intriguing to establish a general proof of the validity of some form of these inequalities for rotating black holes that satisfy the Einstein equations and the dominant energy condition. Alternatively, it would be valuable to identify counterexamples where one or the other of these proposed inequalities is violated.
Furthermore, the bound (48) can be improved if we additionally assume that $p_{r}$ is such that $|r^{3}p_{r}|$ decreases monotonically. In that case, it was proved by Hod Hod (2020) that $r_{\gamma}>\frac{3}{2}r_{H}$, leading to
$\lambda\leq\frac{\sqrt{\mu}}{r}\bigg{|}_{r_{\gamma}}\leq\frac{1}{\sqrt{3}r_{\gamma}}\leq\frac{2}{3\sqrt{3}r_{H}}.$ (49)
Thus, for this family of black holes we find the universal upper bound $\lambda r_{H}\leq\frac{2}{3\sqrt{3}}$. This inequality also implies that for this class the relation
$\lambda^{2}\leq\frac{4\pi}{27S},$ (50)
is satisfied, which is sharper than the bound conjectured in Giataganas et al. (2024) and also to eq.(48). The inequality (50) is saturated by a Schwarzschild black hole.
IV Summary
We have proven that for every static and spherically symmetric black hole satisfying the Einstein equations and the dominant energy condition, there exists a set of universal bounds for the Lyapunov exponent $\lambda$ associated to unstable circular null geodesics in terms of the generalized surface gravity $\kappa_{\gamma}$ Eq.(39) and (40); the Unruh temperature measured by static observers at $r_{\gamma}$, Eq.(42); and the shadow size $R_{sh}$ Eq.(44). Furthermore, we have reformulated some of these bounds in terms of the black hole’s entropy Eq.(48).
For specific black holes where $|r^{3}p_{r}|$ is bounded, we derived a sharper inequality Eq.(50). As a simple application, we show that these bounds also impose a straightforward constraint on the imaginary part of the quasinormal mode frequencies in the eikonal limit.
It would be interesting to know whether some of the bounds proven here have analogs in other spherically symmetric black holes that do not satisfy the assumptions made in this work, specifically regarding the energy conditions or the requirement of being solutions to the Einstein equations. For instance, a generalized version of (19) was proven in for the $n$-dimensional Einstein-Gauss-Bonnet theory, and in Cvetic et al. (2016) Cvetic, Gibbons and Pope have shown that black holes arising from other gravitational theories also satisfy the bounds found inGallo and Villanueva (2015). Since the Lyapunov bound is, in some sense, related to the upper bound on the location of the photon sphere, it is hoped that more general versions of the bound for $\lambda$ exist for broader classes of metrics or alternative theories. More importantly, it would be desirable to test these inequalities for rotating black holes or to prove their validity in those spacetimes. However, in general, the Einstein equations for rotating spacetimes are more complex, and the null geodesics are typically non-planar. Therefore, further work and potentially different techniques will be required to address this problem.
Acknowledgments
E.G acknowledges financial support from CONICET, SeCyT-UNC. E.G. thanks for the motivating discussions to Frisaldo Hames, Dario Gallardo Mozart, and Guillermo Correa Diaz, while T.M. appreciates discussions with Olivera Mišković, Radouane Gannouji, and Rodrigo Olea.